OpenStudy (crazypizzalover):

will fan and medal! What is the solution to the following equation x^2+7x-18=0

OpenStudy (sshayer):

1*-18=-18 9-2=7 9*-2=-18 write 7 as (9-2) and then make factors.

OpenStudy (crazypizzalover):

thank you, what do you mean in make factors?

OpenStudy (crazypizzalover):

@sshayer

OpenStudy (sshayer):

\[x^2+(9-2)x-18=0\] \[x^2+9x-2x-18=0\] \[x(x+9)-2(x+9)=0\] \[(x+9)(x-2)=0\] ?

OpenStudy (crazypizzalover):

I figured it out already thank you so much though!